sir pls give answer of this question
If   x+1/x=3    where x is not equal to 0 then the value of
x4-3x3-5x2-3x-1 / x4-1   is ??
options 
a)5 
b)7
C)2
d)3

Grade:12th pass

1 Answers

Asim
8 years ago
x+ 1 = 3x
 x =( 3 \pm \sqrt{5}) / 2         put these values and check... or … x ^2  +  1/x^2 = 1   thats a hint, check it out its gonna help you a lot... TIY